maths89898:

True or False? k = 3 over 4 is a solution to the inequality 12k + 2 < 12.

Vocaloid:

Thoughts? When you plug in k = 3/4 into 12k + 2, is the result less than 12?

maths89898:

Well let me test

maths89898:

So if k= 3/4

maths89898:

You'd replace k with 3/4 right

Vocaloid:

Yup

maths89898:

So 12k (3/4) + 2 is the equation?

Vocaloid:

Hmm not quite, remember 3/4 replaces k so it’s just 12(3/4) + 2 Keep simplifying from there

maths89898:

Oh right my bad

maths89898:

So what I got is 9x3/4 = 9 + 2 = 11

Vocaloid:

Perfect And since 11 < 12 the inequality is true and k is a solution

maths89898:

Thanks for the help vocaloid!